OpenStudy (dtan5457):

How can we use the partition generating function to solve a problem like this?

OpenStudy (dtan5457):

TO show the number of ways a certain amount of cents can be formed using pennies, nicklels, dimes, and quarters

OpenStudy (dtan5457):

the hard part about this seems to be the limit of integers (only being able to use 1, 5, 10, and 25)

OpenStudy (dtan5457):

@DanJS @Kainui @rational @Blacksteel

OpenStudy (b87lar):

In my understanding, the general idea is as follows: one writes down the generating polynomial that includes all of the listed amounts of cents as powers: \[g(x)=x+x^5+x^{10}+x^{25}\]

OpenStudy (b87lar):

g(x) represents "a coin" Now if someone asks you take k coins and count the number of ways one could get, say, m cents using these k coins, you apply the trick of generating functions. You mutliply k functions g(x): \[g^k(x)=(x+x^5+x^{10}+x^{25})^k\]. You can now multiply this out, perhaps use multinomial fomula to do so. But let us take a concrete example of k=3 coins, and m=25 cents: \[g^3(x)=(x+x^5+x^{10}+x^{25})^3\]

OpenStudy (b87lar):

\[g^3(x)=x^{75}+3 x^{60}+3 x^{55}+3 x^{51}+3 x^{45}+6 x^{40}+6 x^{36}+3 x^{35}+6 x^{31}+x^{30}+3 x^{27}+3 x^{25}+3 x^{21}+3 x^{20}+6 x^{16}+x^{15}+3 x^{12}+3 x^{11}+3 x^7+x^3\]

OpenStudy (dtan5457):

is there for a specific amount of cents though? Like If I asked for the partitions of 26 cents, i don't think there is a restraint on the amount of coins

OpenStudy (b87lar):

(hmm, the line shows partially truncated, but maybe others can see it full). So let me say we are looking for m=40 cents (because that is what seems to fit the display). To get the number of ways three coins can be combined into a sum of 40 cents ais now the factor in front of x^{40}, which turns out to be 6. Six ways.

OpenStudy (b87lar):

you mean given a desired total amount of cents what is the way to get it using any number of coins? I would think the answer is similar, but you need to add the term x^0 to g(x) to account for "no contribution from a particular coin), like this

OpenStudy (b87lar):

\[g(x)=1+x+x^5+x^{10}+x^{25}\]

OpenStudy (b87lar):

then you need to choose k at least that of the desired total amount to account for polynomial to attain the sum using minimal denomitation (1 cent), so that would be g^{40}(x) for 40 cents.

OpenStudy (b87lar):

Generally you want to find some formula, similar to the multinomial formula, that gives you a prescription for the factor of each term, so you do not really need to multiply everything out. Does this make sense?

OpenStudy (dtan5457):

im still a little confused.

OpenStudy (dtan5457):

possible you can demonstrate a full out example for any number of cents?

OpenStudy (b87lar):

when you say "any number of cents" are you referring to the total sum (value in cents) or to the number of coins?

OpenStudy (dtan5457):

i would like it to be the total sum. but if it's a lot more complicated then we can stick with the number of coins

OpenStudy (b87lar):

okay we can take an example: 40 cents using any number of coins

OpenStudy (dtan5457):

ok

OpenStudy (b87lar):

so i took the 40th power of the polynomial g(x) = 1+x+x^5+x^10+x^25. The coefficient of the power 40 of that polynomial is astronomical, about 7000 trillion.

OpenStudy (b87lar):

which makes sense if one thinks about having 40 1 cent coins permuted all ways, then 35 pennies permuted with 1 nickel, etc.etc.

OpenStudy (b87lar):

Normally, the number of coins is given which makes the problem manageable, like I showed further above.

OpenStudy (dtan5457):

i was hoping for this to be an ordered partition is that possible

OpenStudy (b87lar):

the main point is that you get the essence of why this works, then you will solve any similar problem

OpenStudy (b87lar):

what precisely do you mean by ordered partition?

OpenStudy (dtan5457):

if it was ordered wouldnt it be alot more manageable? like you can only arrange 40 pennies once.

OpenStudy (dtan5457):

thats actually called non-ordered my bad

OpenStudy (dtan5457):

if we allowed order partitions that formula is generally 2^(x-1)

OpenStudy (dtan5457):

then generating is usually for non-ordered

OpenStudy (b87lar):

actually my comment on permutation was out of place - the arragement of 40 pennies will be counted only once in the total number of ways we can get 40 cents.

OpenStudy (dtan5457):

right

OpenStudy (b87lar):

so let me make sure i understand correctly: for a simpler example how many non-ordered arrangements of pennies and dimes can get you total of 6 cents,

OpenStudy (b87lar):

(sorry i mean to say nickel instead of dime) you want to count 6 pennies as one, then 1 nickel + 5 pennies and that's it, i.e. 2 arrangements?

OpenStudy (dtan5457):

sounds right

OpenStudy (dtan5457):

obviously dimes and quarters can be used too but they cant in that situation

OpenStudy (dtan5457):

but if its near impossible. let's just stick with your first idea.

OpenStudy (b87lar):

i need to sign off but will keep thinking about this. thanks

OpenStudy (dtan5457):

no, thank you :) i still have a day to figure this out thanks for taking your time

OpenStudy (b87lar):

This seems to have a bunch of useful info and might actually answer your question https://math.berkeley.edu/~mhaiman/math172-spring10/partitions.pdf. I haven't had time to read through but will do so, too.

OpenStudy (dtan5457):

im reading it and im looking for the part that will specificlaly help me, know where that is?

OpenStudy (dtan5457):

@satellite73

OpenStudy (dtan5457):

but im actually getting more and more interested in your first interpretation of the problem as well

OpenStudy (dtan5457):

would it be (x+x^5+x^10+x^25)^k for any number of cents?

OpenStudy (b87lar):

yes, with k coins

OpenStudy (dtan5457):

oh no wait i think i know what you mean. you put k as the amount of coins and all the exponents are the possible cents u can get with 3 coins, thats why so many exponents are missing because alot of number of cents cannot be made with only 3 coins, correct?

OpenStudy (b87lar):

you got it!

OpenStudy (dtan5457):

im definitely gonna use that for my research paper. however for the final one due in a week, if its possible to solve my interpretation of the problem, that would be nice as well but thanks anyhow

OpenStudy (dtan5457):

also, do you know how to do that with non-ordered partitions?